sqrt3 sqrt3
----- - -----
sqrt6 -1 sqrt6 + 1

sqrt9x^2 = x + 6
I am having a hard time figuring these out. Can you help?(sqrt means square root)
: sqrt3 sqrt3
----- - -----
sqrt6 -1 sqrt6 + 1

sqrt9x^2 = x + 6
I am having a hard time figuring these out. Can you help?(sqrt means square root)

Answer by stanbon(18787)

sqrt3 sqrt3
----- - -----
sqrt6 -1 sqrt6 + 1
------------------------
To subtract fractions you need to have a common denominator on the fractions.
-------------
Least Common Denominator: (sqrt(6)-1)(sqrt(6)+1) = 6-1 = 5
------------------------
Rewrite the problem with lcm = 5
[(sqrt(3)(sqrt(6)+1) - sqrt(3)(sqrt(6)-1)]/5
[(sqrt(18) + sqrt(3)) - (sqrt(18)-sqrt(3))]/5
= [2sqrt(3)] / 5
= (2/5)*sqrt(3)
=====================
Cheers,
Stan H.